OpenStudy (anonymous):

x^4-16^2+15=0 please solve

OpenStudy (angela210793):

x^2-241=0 x^2=241 x=+-sqr241

OpenStudy (diyadiya):

\[x=\sqrt[4]{271}\]

OpenStudy (anonymous):

i'm sorry! i meant x^4-16x^2+15=0

OpenStudy (anonymous):

4.0573485964548730093221785586954571686819749139665696236896

OpenStudy (anonymous):

please factor :D

OpenStudy (anonymous):

(x-1) (x+1) (x^2-15) = 0

OpenStudy (anonymous):

sorry but could you explain how you got that?

OpenStudy (anonymous):

x^4 - 16x^2 + 15=0 possible factors of x^4: x^2 &x^2 possible factors of 15: -15 & -1 cross multiply and add = -16x^2 (middle term) -15